OpenStudy (garretttyre05):

Plz i need help asap. Will fan and medal. Find a rational number and an irrational number that are between 5.2 and 5.5. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (welshfella):

Hint a ration number can be written as a fraction so 5 plus any fraction between 1/5 and 1/2 would be a rational number.

OpenStudy (welshfella):

a rational number would be a surd like sqrt20 for example so you need to think af a number between 5.2^2 and 5.5^2 and take the square root of it.

OpenStudy (welshfella):

* correction thats an irrational number

OpenStudy (welshfella):

can you think of a fraction between 1/5 and 1/2? Or a decimal between 5.2 and 5.5 ( thats the easiest way).

OpenStudy (garretttyre05):

So i can put 5.3 and that would be right?

OpenStudy (welshfella):

Yes thats great

OpenStudy (garretttyre05):

But what about the irrational number?

OpenStudy (welshfella):

5.2^2 = 27.04 5.5^2 = 30.25 - pick any whole number between these 2 which is not a perfect square

OpenStudy (garretttyre05):

27.04

OpenStudy (garretttyre05):

then would i just keep repeating the 04

OpenStudy (welshfella):

no a whole number between this and 30.25

OpenStudy (garretttyre05):

so 29

OpenStudy (welshfella):

yea thats fine because its not a square number so the irratuional number we want is sqrt 29 If you grab your calculatopr you'll see that will have a value of just below 5.5

OpenStudy (garretttyre05):

and have a dumb question what the decimal approximation

OpenStudy (welshfella):

5.385164807

OpenStudy (garretttyre05):

whats*

OpenStudy (garretttyre05):

ok so for the rational number the answer is 5.3 and the answer for the irrational number is 29?

OpenStudy (welshfella):

well you cant write an irrational number accurately as a decimal - the numbers go on infinitely. Most calculators e you up to 8 decimal places.

OpenStudy (garretttyre05):

so then whats the answer?

OpenStudy (welshfella):

they want it to nearest hundredth Thats the second place of decimals 5.385164807 to nnearesrt hundredth is 5.39 29 is not the answer The answer is sqrt 29 and 5.39 is an approximation of sqrt 29.

OpenStudy (welshfella):

do you follow the reasoning ok?

OpenStudy (welshfella):

- remember tha 2 numbers had to be beween 5.2 and 5.5.

OpenStudy (garretttyre05):

ok so is that the answer for rational or irrational?

OpenStudy (welshfella):

rational is 5.3 and irrational is sqrt29 or 5.39 to nearest hundredth

OpenStudy (garretttyre05):

ok rational is 5.3 and irrational is \[\sqrt{5.39}\]

OpenStudy (welshfella):

no irrational is \[\sqrt{29}\]

OpenStudy (garretttyre05):

ok i got it know

OpenStudy (welshfella):

but they required a decimal approximation of the irrational number - that is 5.39

OpenStudy (garretttyre05):

so i would just put 5.39

OpenStudy (welshfella):

so ther are 3 answers 5.3 , sqrt 29 and 5.39

OpenStudy (garretttyre05):

but the one i need to put is 5.39

OpenStudy (welshfella):

the first part of the question wanted a rational and irrational number between 5.2 and 5.5. The answer to that is 5.3 and \[\sqrt{29}\]

OpenStudy (welshfella):

The last part wanted the irational number as a decimal to nearest hundedrth that answer to that part is 5.39.

OpenStudy (welshfella):

ok??